V, Therefore the l a center at (2,-3), and a s shown at right. Foblems For each of the following equations, state the parent equation and then sketch its graph. Be sure to include any key and/or locator points. 1. (x-5)2 У (x-2)2 + (y +1)2-9 y 1 + 10 y -(x-2)3 + 1 4. y +5-2 6. y 2-8 5. For each of the following problems, state whether or not it is a f explain why not.Can someone show me how to do number 5 please? I'm really struggling on that one
Answer #1

5)

We will use the parent function y =1/x to sketch the graph of y= 1/(x+1) +10 using transformation.

Consider the well known Graph of y=1/x

4 2 2 4

Now we will shift the above graph 1 unit to the left and 10 units UP to obtain the graph of y= 1/(x+1) +10